OpenStudy (nathanjhw):

The expression dy/dx = x(3√y) gives the slope at any point on the graph of the function f(x) where f(2) = 8.

OpenStudy (nathanjhw):

A. Write the equation of the tangent line to f(x) at point (2, 8). B. Write an expression for f(x) in terms of x. C.What is the domain of f(x)? D. What is the minimum value of f(x)? E. Using the axes provided, sketch a slope field for the given differential equation at the nine points indicated.

OpenStudy (nathanjhw):

@welshfella

OpenStudy (nathanjhw):

@phi

OpenStudy (welshfella):

A:- substitute the values of m ( = dy/dx) and the point (2,8) in the standard form y - y1 = m(x - x1) x1 = 2 and y1 = 8 NOte also :- x = 2 and y =8 for dy/dx

OpenStudy (nathanjhw):

So are you saying the equation is y-8= m(x-2)?

OpenStudy (nathanjhw):

@welshfella

OpenStudy (welshfella):

yes but you can work out the slope m from the info given m = dy/dx

OpenStudy (nathanjhw):

m=8/2 = 4?

OpenStudy (welshfella):

plug in x =2 and y = 8 into dy/dx

OpenStudy (welshfella):

dy/dx = 2(3sqrt8)

OpenStudy (welshfella):

that = m in the dtandard equation

OpenStudy (nathanjhw):

so we get y-8 = 2(3sqrt8)(x-2)

OpenStudy (welshfella):

yes but we could simplify it a bit y = 6sqrt8 (x - 2) + 8

OpenStudy (nathanjhw):

Alright that makes sense.

OpenStudy (welshfella):

sorry gotta go right now

OpenStudy (nathanjhw):

That's fine. You helped me with the first part.

OpenStudy (nathanjhw):

@rational can you pick up where he left off?

OpenStudy (nathanjhw):

@jigglypuff314 could you pickup where he left off?